OpenStudy (mendicant_bias):

Trying to solve another induction proof problem (one part of a larger problem) and just trying to remember how to properly do this. Prompt and rough work below.

OpenStudy (mendicant_bias):

\[\text{Given} \ x_{n+1} = \frac{1+x_n}{2}; \ \ \ x_1 > 1,\]\[\text{a.) Use Induction to prove} \ (\forall n \in N), \ \ \ x_n >1.\]

OpenStudy (mendicant_bias):

@ganeshie8

OpenStudy (mendicant_bias):

I never really got the hand of answering induction proofs that weren't built around a series but rather an inequality or equality, how do I do this?

OpenStudy (mendicant_bias):

I guess I also don't understand on the inductive step how we know which equation to modify or use as the "assume x_k ..." relation. Which do we pick?

OpenStudy (mendicant_bias):

And I know the first step, it's just given in this problem to my understanding, x_1 > 1, right.

ganeshie8 (ganeshie8):

Hey still here ?

OpenStudy (mendicant_bias):

I'm here now.

ganeshie8 (ganeshie8):

Here is the complete proof : Base case : \(x_1\gt 1\) Induction Hypothesis : Assume \(x_{k} \gt 1\) Induction step : \(x_{k+1} = \dfrac{1}{2}+\dfrac{1}{2}*x_k\gt \dfrac{1}{2}+\dfrac{1}{2}*1=1~~\blacksquare\)

OpenStudy (mendicant_bias):

Ah, alright. Thank you!